Background Information:

One of the most common forms of protection programs in use today are retirement programs. While there are several different types of retirement programs available to employees in various industries, one of the most common and well-known retirement programs is the 401(k) plan. For this homework, you will be practicing calculations that help you explore both the 401(k) contribution and employer matching process.

Instructions:

You have just started a new job and are thrilled to learn that your new employer offers a 401(k) retirement plan to its employees. Your annual salary is $40,000. Assume the IRS allows you to contribute up to $24,000 to your 401(k). You’ve decided to contribute 7% of your annual salary to the plan.

Questions:

1. How much more money would you need to contribute to meet the maximum allowable contribution set forth by the IRS?

2. The company offers you a $.50 match for each dollar that you contribute between 2 and 5 percent of your annual salary. How much is the company match based on your 7% contribution?

3. Is this a defined benefit plan or defined contribution plan? Why?
Answer #1

Question 1

Annual Salary = $40,000

Contribution = 7%

Annual Contribution = $2800

Maximum Contribution allowed = $24,000

Difference = 24000-2800 = $21,200

Question 2

Dollar contribution between 2% and 5%

2% contribution = $800

5% contribution = $2000

Total contribution between 2% and 5% = 2000-800 = $1200

Company's contribution = .5 of 1200 = $600

Question 3

Its a defined contribution plan. this is because the company is committing to contribute a particular sum of money but is not guaranteeing any future benefit.